OpenStudy (anonymous):

Consider the system of equations: x + y + z = 210; x + z = 230; x + y = 120 What is the value of x?

OpenStudy (anonymous):

First set up a system of linear equations. x + z = 230 x + y = 120 ----------- 2x + y + z = 350 Now, set up a new system. 2x + y + z = 350 x + y + z = 210 What can you do to the second equation to eliminate x and y?

OpenStudy (anonymous):

3x+2y+2z=560

OpenStudy (anonymous):

Umm...no. What do you do to the second equation so that WHEN you add them, the y and z's cancel?

OpenStudy (anonymous):

im not sure boss i told my instructor i was having problems with these and she just gave me websites to look at

OpenStudy (anonymous):

2x + y + z = 350 x + y + z = 210 What happens when you multiply the entire second equation by -1?

OpenStudy (anonymous):

it just negates itself

OpenStudy (anonymous):

Yeah. 2x + y + z = 350 -x - y - z = -210 Now add the equations together and see what happens.

OpenStudy (anonymous):

x=140

OpenStudy (anonymous):

Yup. Look at the question. It asks for the value of x. So, there you have it. Do you understand why we multiplied the second equation by -1 though?

OpenStudy (anonymous):

to negate or get rid of the y and z

OpenStudy (anonymous):

You multiply by -1 so that the y and z cancel out when you add them. THis is a process called elimination.

OpenStudy (anonymous):

you help out tremendously and explained so i can understand thank you so much

OpenStudy (anonymous):

so it ends up 140+-20+90=210 and checks out

OpenStudy (campbell_st):

given x + y + z = 210 and x+ z = 230 then (x + z) + y = 210 or 230 + y = 210 y = -20 and x + y = 120 then x - 20 = 120 x = 140

OpenStudy (anonymous):

Yeah. What @campbell_st did is something called substitution.

OpenStudy (anonymous):

thanks fellas

OpenStudy (anonymous):

np :)